OpenStudy (anonymous):

Divide: 8-3i div5-4i

OpenStudy (anonymous):

\[\frac{8-3i}{5-4i}\times\frac{5+4i}{5+4i}\]

OpenStudy (anonymous):

(8-3i)/(5-4i) =[(8-3i)(5+4i)]/[(5-4i)(5+4i)]=(52+17i)/(5^2+4^2) =(52+17i)/(41)

OpenStudy (anonymous):

\[\frac{40+17i-12i^2}{25-16i^2}= \frac{40+17i-12(-1)}{25-16(-1)}\]\[\frac{40+12+17i}{25+16}=\frac{52+17i}{41}\]